# Magneto-optical properties of Group-IV–vacancy centers in diamond upon hydrostatic pressure

### v1

@@ III Results
In this study, we characterize the G4V(--) color centers under compressive hydrostatic pressure by means of ab initio supercell calculations. In this context, we always mean compressive pressure in the entire study that we do not explicitly mention further on. Under hydrostatic pressure, the fundamental band gap of the diamond widens that can also shift the photoionization threshold levels of G4V defects which sets the photoionization thresholds of G4V(--) color centers. So, we compute first the charge transition levels of G4V defects to identify their photostable excitation energies assuming single photon absorption. Then we compute the ZPL energies under increasing hydrostatic pressure up to 180 GPa which gives information about the conditions of fluorescence for each G4V(--) color center at a given hydrostatic pressure. We continue the characterization of the fine level structure in the electronic ground and excited states under hydrostatic pressure. In particular, the spin-orbit splitting is determined together with hyperfine tensors for the relevant isotopes of the dopant and neighbor 13C nuclear spins as well as the quadrupole moments of the relevant dopants in G4V(--) centers. These data are spectroscopy fingerprints of the G4V(--) centers under a given hydrostatic pressure. Furthermore, these data are also important in understanding the temperature-dependent electron spin coherence time of G4V(--) color centers. These spin interactions are strongly interwoven with electron-phonon interaction that we treat with Jahn–Teller theorem as we briefly described in the Method section. As a consequence, the Jahn–Teller parameters are calculated at each considered hydrostatic pressure for each G4V(--) color center at the electronic ground and excited states and will be reported before providing the fine level structure.
@@ III.1 Photoionization thresholds and zero-phonon line energies
Figure 2: (a) Calculated charge transition levels of SiV, GeV, SnV and PbV defects under hydrostatic pressure in the range of 00 to $180$ GPa as obtained by HSE06 functional. (b) The ZPL shift under hydrostatic pressure was calculated by SCAN functional where the zero pressure value was aligned to the experimental data. For SiV and GeV defects the experimental data points for non-zero pressure are also given from Ref. Vindolet et al., 2022. We highlight the hydrostatic pressure region by light red color where PbV(--) is not photostable.
@@
The photoionization threshold energies can be calculated from the charge transition levels referenced to the appropriate band edge. In particular, the photostability of G4V(--) is of interest. In this case, the (|0)(-|0)( - | 0 ) level with respect to the conduction band minimum (CBM) yields the photoionization threshold for converting the negatively charged defect to neutral one, whereas the (2|)(2-|-)( 2 - | - ) level with respect from the valence band maximum (VBM) yields the photoionization threshold energy for converting the negatively charged defect to doubly negative charged one. If the ZPL energy of the G4V(--) is larger than either of these photoionization threshold energies then G4V(--) is not a photostable color center.
@@
The charge transition levels of defects can be calculated within $0.1$ eV accuracy in diamond by HSE06 functional Deák et al. (2010), thus we apply this method. The calculation methods of charge transition levels for G4V defects are given in our previous study Thiering and Gali (2018) with using the standard procedure Freysoldt et al. (2014) that we do not reiterate here. We note that the reported values are valid at $T=0$ K and phonon-assisted photoionization can decrease the effective thresholds at elevated temperatures (e.g., Ref. Csóré et al. (2022)).
@@
We find that the increase of hydrostatic pressure from 00 to $180$ GPa leads to the widening of the fundamental band gap of diamond by $0.66$ eV. The calculated charge transition levels of G4V(--) shift upward with increasing pressure. The general trend is that the heavier the dopant ion the steeper is the upward shift in the charge transition level (Fig. 2). The same trend can be observed in the shift of ZPL. We already discussed its origin in our previous study for SiV(--) and GeV(--) color centers Vindolet et al. (2022) which can be qualitatively explained through an investigation of Kohn–Sham molecular orbital states by using a single-particle picture. In the first order, the ZPL energy scales with the difference between the unoccupied and occupied states, named $e_{g}$ and $e_{u}$. Thus, the change in ZPL energy upon varying hydrostatic pressures can be estimated as $E_{\text{ZPL}}(p)\propto\varepsilon_{u}(p)-\varepsilon_{g}(p)$, where $\varepsilon_{g}(p)$ and $\varepsilon_{u}(p)$ are the Kohn–Sham energies of the orbitals $e_{g}$ and $e_{u}$ at pressure $p$, respectively. The crucial distinction among the hydrostatic pressures is the deformation potential of the $e_{g}$ orbital, which transitions from bonding to antibonding as the pressure increases Harrison (2012). This applies to SnV(--) and PbV(--) color centers too. Although PbV(--) shows the most sensitive pressure dependence, this also holds to the photoionization thresholds. We find that the photoionization threshold towards CBM crosses the ZPL energies at around $32$ GPa hydrostatic pressure which sets a limit for PbV(--) in pressure sensor application.
@@ III.2 Dynamic Jahn-Teller effect and electron-phonon parameters
The electronic structure of the electronic ground and excited states is subject to the dynamic Jahn–Teller (DJT) effect for G4V(--) centers that was described in detail in Ref. Thiering and Gali, 2018 and later exploited in other studies too Krivobok et al. (2020); Vindolet et al. (2022); Harris et al. (2023). In DFT calculations, the geometry optimization yields the low $C_{2h}$ symmetry as the global energy minimum for zero pressure but the system can be forced to stay in the high $D_{3d}$ symmetry in the DFT calculations. In $D_{3d}$ symmetry the distances between the three symmetrically equivalent neighbor carbon atoms are labeled by $d_{2}$ whereas the distance between the dopant ion and the nearest neighbor carbon atoms is labeled by $d_{1}$ in Fig. 1(a), and the values are listed in Table 1. The general trend is that $d_{1}$ and $d_{2}$ values increase with heavier ions in G4V(--) centers at zero pressure. This may be expected as a larger ion creates a stronger strain field around the dopant ion. For a given G4V(--) center the $d_{1}$ and $d_{2}$ values slowly decrease with increasing the hydrostatic pressure. This may be also expected as the hydrostatic pressure generally decrease the bond length in the diamond crystal which also holds for the bond lengths in the center of G4V defects.
@@
SiV(--) 1.97///1.96111Ref. Thiering and Gali, 2018 2.67///2.67111Ref. Thiering and Gali, 2018 1.79 2.40
GeV(--) 2.01///2.01111Ref. Thiering and Gali, 2018 2.74///2.73111Ref. Thiering and Gali, 2018 1.84 2.48
SnV(--) 2.09///2.08111Ref. Thiering and Gali, 2018 2.83///2.83111Ref. Thiering and Gali, 2018 1.92 2.58
PbV(--) 2.12///2.12111Ref. Thiering and Gali, 2018 2.89///2.88111Ref. Thiering and Gali, 2018 1.96 2.64
@@
Table 2: The calculated Jahn-Teller (JT) energy in meV ($E_{\text{JT}}$), the barrier energy in meV ($\delta_{\text{JT}}$), the effective mode in meV ($\hbar\omega$), spin-orbit splitting $\lambda_{0}$ (meV), $p$ reduction factor, and $\lambda$ (GHz) which is Ham reduced $\lambda_{0}$ for the G4V(--) defects ($\lambda=p\lambda_{0}$) with no applied pressure. The left/right hand side data are from SCAN/HSE06 calculations.
Defects $E_{\text{JT}}$ (meV) $\delta_{\text{JT}}$ (meV) $\hbar\omega$ (meV) $\lambda_{\text{0}}$ (meV) $p$ $\lambda$ (GHz) $\lambda_{\text{exp}}$ (GHz)
@@
SiV(--) 40.86 / 42.30111Ref. Thiering and Gali, 2018 3.79 / 3.0111Ref. Thiering and Gali, 2018 89.70 / 85.20111Ref. Thiering and Gali, 2018 0.86 / 0.82111Ref. Thiering and Gali, 2018 0.34 / 0.31111Ref. Thiering and Gali, 2018 70.26 / 61.0111Ref. Thiering and Gali, 2018 50222Ref. Hepp et al., 2014
GeV(--) 30.59 / 30.10111Ref. Thiering and Gali, 2018 4.05 / 2.0111Ref. Thiering and Gali, 2018 77.01 / 82.20111Ref. Thiering and Gali, 2018 2.45 / 2.20111Ref. Thiering and Gali, 2018 0.38 / 0.39111Ref. Thiering and Gali, 2018 222.8 / 207111Ref. Thiering and Gali, 2018 181333Ref. Ekimov et al., 2015
SnV(--) 20.81 / 21.60111Ref. Thiering and Gali, 2018 1.15 / 1.6111Ref. Thiering and Gali, 2018 64.87 / 79.40111Ref. Thiering and Gali, 2018 8.69 / 8.28111Ref. Thiering and Gali, 2018 0.44 / 0.47111Ref. Thiering and Gali, 2018 915.2 / 946111Ref. Thiering and Gali, 2018 850444Ref. Iwasaki et al., 2017b
PbV(--) 15.02 / 15.60111Ref. Thiering and Gali, 2018 3.88 / 0.6111Ref. Thiering and Gali, 2018 51.98 / 74.90111Ref. Thiering and Gali, 2018 35.0 / 34.6111Ref. Thiering and Gali, 2018 0.48 / 0.54111Ref. Thiering and Gali, 2018 4097 / 4514111Ref. Thiering and Gali, 2018 3914555Ref. Wang et al., 2021
@@
SiV(--) 62.58 / 78.50111Ref. Thiering and Gali, 2018 1.12 / 2.7111Ref. Thiering and Gali, 2018 60.97 / 73.50111Ref. Thiering and Gali, 2018 8.864 / 6.96111Ref. Thiering and Gali, 2018 0.133 / 0.128111Ref. Thiering and Gali, 2018 286.2 / 215111Ref. Thiering and Gali, 2018 260222Ref. Hepp et al., 2014
GeV(--) 71.48 / 85.70111Ref. Thiering and Gali, 2018 2.31 / 5.4111Ref. Thiering and Gali, 2018 70.64 / 73.00111Ref. Thiering and Gali, 2018 35.03 / 36.1111Ref. Thiering and Gali, 2018 0.136 / 0.113111Ref. Thiering and Gali, 2018 1155 / 987111Ref. Thiering and Gali, 2018 1120333Ref. Ekimov et al., 2015
SnV(--) 67.69 / 83.10111Ref. Thiering and Gali, 2018 4.20 / 6.8111Ref. Thiering and Gali, 2018 68.13 / 75.60111Ref. Thiering and Gali, 2018 94.77 / 96.8111Ref. Thiering and Gali, 2018 0.140 / 0.125111Ref. Thiering and Gali, 2018 3214 / 2897111Ref. Thiering and Gali, 2018 3000444Ref. Iwasaki et al., 2017b
PbV(--) 87.32 / 91.60111Ref. Thiering and Gali, 2018 6.69 / 12.3111Ref. Thiering and Gali, 2018 77.93 / 78.60111Ref. Thiering and Gali, 2018 241.4 / 245111Ref. Thiering and Gali, 2018 0.116 / 0.119111Ref. Thiering and Gali, 2018 6782 / 7051111Ref. Thiering and Gali, 2018 -
For each applied hydrostatic pressure, we calculated the APES by SCAN functional as illustrated in Fig. 1(d). For the zero-pressure case, we compared the SCAN data with previous HSE06 data (see Table 2). We find a good agreement between the respective SCAN data and HSE06 data for the critical JT parameters (such as the Jahn–Teller energy $E_{\text{JT}}$, barrier energy $\delta_{\text{JT}}$, and the effective frequency $\hbar\omega$) so we continued the calculations with SCAN functional for non-zero pressure cases. We observe the same DJT effect under hydrostatic pressure as for the no-pressure case with resulting in $C_{2h}$ geometry in the global energy minimum. We list the respective JT parameters for G4V(--) centers under pressure in Table 4 in the Appendix.
@@ III.3 Effective spin-orbit splitting
We calculated the spin-orbit coupling as obtained by SCAN functional which is an essential parameter to compute the fine electronic structure of G4V(--) centers. We label this electronic spin-orbit coupling as $\lambda_{0}$ and is listed for each G4V(--) for the zero-pressure case in Table 2. We find that the SCAN data and the previously published HSE06 $\lambda_{0}$ values are in good agreement for each electronic state in each G4V(--) color center. Therefore, we are convinced that the pure electronic spin-orbit coupling can be further calculated for the G4V(--) color centers under non-zero hydrostatic pressures with SCAN functional.
@@
Figure 3: (a) The SCAN calculation of hydrostatic pressure (GPa) dependence of the effective spin-orbit splitting $\lambda$ in GHz unit (GHz) for G4V(--) color centers in the electronic excited and ground states plotted by blue and red lines, respectively. (b) The sum of the electronic excited and ground state’s $\lambda$ values associated with the zero-phonon line broadening at elevated observation temperatures with off-resonant excitation for each G4V(--) color center as a function of the applied hydrostatic pressure.
@@
The electronic spin-orbit coupling is reduced due to strong electron-phonon coupling, which is often called the Ham reduction factor that we label by $p$. We described the ab initio calculation of the Ham reduction factor for the relevant $E\otimes e$ JT system in our previous publication Thiering and Gali (2017) that we applied to G4V(--) defects in a subsequent study Thiering and Gali (2018). We find for the non-zero pressure case that the calculated $p$ factors and the resulting $\lambda$ for the electronic ground and excited state of G4V(--) as obtained by SCAN and HSE06 functionals agree and they also agree with the experimental data (see Table 2) as expected because both the JT parameters and the electronic spin-orbit values agree well as obtained by the respective SCAN and HSE06 functionals. We conclude that we can further apply the SCAN functional to study the hydrostatic pressure dependence of the effective spin-orbit splitting in the G4V(--) color centers.
@@
The calculated effective spin-orbit slitting under hydrostatic pressure as obtained by SCAN functional is plotted in Fig. 3(a). The general trend is that the effective spin-orbit splitting is monotonously increasing with increasing hydrostatic pressure both in the electronic ground and excited states. The calculated effective spin-orbit splitting has two factors: the Ham-reduction parameter and the calculated electronic spin-orbit splitting (see Table 4). We find that the Ham-reduction parameter only slightly varies as a function of hydrostatic pressure, thus the inherent electronic spin-orbit splitting is dominantly increasing with raising the hydrostatic pressure. The largest contribution to the electronic spin-orbit splitting comes from the spin-orbit coupling on the dopant ion Thiering and Gali (2018). With increasing the hydrostatic pressure the distance ($d_{1}$) between the dopant ion and the neighbor carbon atoms is decreasing so the overlap of the dangling bonds with the dopant ion is increasing (see also Sec. III.4). As a consequence, the wavefunction cloud around the dopant ion is denser with increasing hydrostatic pressure which contributes the most to the increasing spin-orbit coupling. The steepest shift in the effective spin-orbit splitting can be observed for SnV(--) color center.
@@
We plot the sum of $\lambda_{g}$ and $\lambda_{u}$ defined in Fig. 1(c) - as a function of the applied hydrostatic pressure in Fig. 3 that we discuss in the context of the broadening of the ZPL emission. Off-resonant excitation of G4V(--) color centers should result in emission in all the possible combinations of states and levels depicted in Fig. 1(c). By assuming a homogeneous strain field and neglecting the isotope effects (in the GHz region), the sum of $\lambda_{g}$ and $\lambda_{u}$ yields the width of the zero-phonon line. In experiments, the broadening of ZPL for SiV(--) varied from 10 meV to 30 meV going from zero to $180$ GPa of hydrostatic pressure, whereas it varied from 15 meV to 30 meV for GeV(--), respectively Vindolet et al. (2022). In our calculations, the spin-orbit splitting related broadening of the ZPL emission varies between 1.4 meV to 1.8 meV for SiV(--) going from zero to $180$ GPa of hydrostatic pressure, whereas it varies between 5.5 meV to 6.7 meV for GeV(--), respectively. The calculated values are an order of magnitude smaller both in terms of absolute values and changes upon hydrostatic pressures. We conclude that before applying pressure to the sample, non-uniform local strain fields could be already present in the samples, which causes inhomogeneous broadening in the region of 10 meV. By applying hydrostatic pressure to the samples, it could further increase the variety of the local strain fields that may explain the observed wide broadening of the ZPL emission for SiV(--) and GeV(--) optical centers.

### v2

@@ III Results
Throughout this study, we characterize the magneto-optical properties of the negatively charged G4V(--) color centers under compressive hydrostatic pressure. While not explicitly stated henceforth, all references to pressure will assume compressive hydrostatic conditions.
Under hydrostatic pressure, the fundamental band gap of the diamond widens, which also shifts the photoionization threshold levels of G4V defects. To account for this effect, we first compute the charge transition levels of G4V defects to identify their photostable excitation energies, assuming single photon absorption. Then, we compute the ZPL energies under increasing hydrostatic pressure up to 180 GPa, which gives information about the fluorescence conditions for each G4V(--) color center at a given hydrostatic pressure.
After exploring charge dynamics under optical illumination, we focus on the effects of compressive hydrostatic pressure on the fine-level structure of G4V(--) centers in the electronic ground and excited states. In particular, the spin-orbit splitting is determined with hyperfine tensors for the relevant isotopes of the dopant and neighbor 13C nuclear spins and the quadrupole moments of the relevant dopants in G4V(--) centers. These data serve as spectroscopic fingerprints of the G4V(--) centers under specific hydrostatic pressure conditions. Furthermore, this information is essential in understanding the temperature-dependent electron spin coherence times of G4V(--) color centers. These spin interactions are strongly interwoven with electron-phonon interaction, which we treat with Jahn–Teller theory, as briefly described in the Method section. Consequently, the Jahn–Teller parameters are calculated at each considered hydrostatic pressure for each G4V(--) color center at the electronic ground and excited states and will be reported before providing the fine-level structure.
@@ III.1 Photoionization thresholds and zero-phonon line energies
Figure 2: (a) Calculated charge transition levels of SiV, GeV, SnV and PbV defects under hydrostatic pressure in the range of 00 to $180$ GPa as obtained by HSE06 functional. (b) The ZPL shift under hydrostatic pressure was calculated by SCAN functional where the zero pressure value was aligned to the experimental data. For SiV and GeV defects the experimental data points for non-zero pressure are also given from Ref. 19. We highlight the hydrostatic pressure region by light red color where PbV(--) is not photostable. We list the charge correction to the total energy and dielectric constant dependency on pressure in Table 8.
@@
The photoionization threshold energies can be calculated from the charge transition levels referenced to the appropriate band edge. In particular, the photostability of G4V(--) is of interest. In this case, the (|0)(-|0)( - | 0 ) level with respect to the conduction band minimum (CBM) yields the photoionization threshold for converting the negatively charged defect to a neutral one, whereas the (2|)(2-|-)( 2 - | - ) level with respect to the valence band maximum (VBM) yields the photoionization threshold energy for converting the negatively charged defect to doubly negative charged one. If the ZPL energy of the G4V(--) center exceeds either of these photoionization threshold energies, the G4V(--) center cannot be considered photostable. In this case, excitation energies at or above the ZPL could induce photoionization, transferring the defect to an alternative charge state.
@@
The charge transition levels of defects in diamond can be accurately calculated using the HSE06 functional, with an accuracy of approximately $0.1$ eV [59]. Therefore, we employ this method in our study. The computational approach for determining the charge transition levels of G4V defects follows the standard procedure outlined in our previous study [26] and is based on the methodology detailed by Freysoldt et al. [60], which we do not reiterate here. We note that the reported values are valid at $T=0$ K, and phonon-assisted photoionization can decrease the effective thresholds at elevated temperatures (e.g., Ref. [61]).
@@
We find that the increase of hydrostatic pressure from 00 to $180$ GPa leads to the widening of the fundamental band gap of diamond by $0.66$ eV. The calculated charge transition levels of G4V(--) shift upward with increasing pressure. The general trend is that the heavier the dopant ion, the steeper the upward shift in the charge transition level (Fig. 2). The same trend can be observed in the shift of ZPL. We already discussed its origin in our previous study for SiV(--) and GeV(--) color centers [19], which can be qualitatively explained through an investigation of Kohn–Sham molecular orbital states using a single-particle picture. In the first order, the ZPL energy scales with the difference between the unoccupied and occupied states, named $e_{g}$ and $e_{u}$. Thus, the change in ZPL energy upon varying hydrostatic pressures can be estimated as $E_{\text{ZPL}}(p)\propto\varepsilon_{u}(p)-\varepsilon_{g}(p)$, where $\varepsilon_{g}(p)$ and $\varepsilon_{u}(p)$ are the Kohn–Sham energies of the orbitals $|e_{g}\rangle$ and $|e_{u}\rangle$ at pressure $p$, respectively. The crucial distinction among the response to hydrostatic pressures is the deformation potential of the $|e_{g}\rangle$ orbital, which transitions from bonding to antibonding character as the atomic number of impurity atom increases [19]. Although PbV(--) shows the most sensitive pressure dependence, this also applies to photoionization thresholds. We find that the photoionization threshold towards CBM crosses the ZPL energies around $32$ GPa hydrostatic pressure, which sets a limit for PbV(--) in pressure sensor applications.
@@ III.2 Dynamic Jahn-Teller effect and electron-phonon parameters
The electronic structure of both the ground and excited states of the G4V(--) centers is subject to the dynamic Jahn–Teller effect, as thoroughly characterized in Ref. 26 and subsequently utilized in further studies [15, 19, 35]. In DFT calculations, geometry optimization follows the sombrero hat potential, resulting in a low-symmetry $C_{2h}$ geometry with the global energy minimum at zero pressure. However, the system can be constrained to retain the high-symmetry $D_{3d}$ configuration by employing the fractional occupations described above. In $D_{3d}$ symmetry, the distances between the three symmetrically equivalent neighbor carbon atoms are labeled by $d_{2}$, whereas the distance between the dopant ion and the nearest neighbor carbon atoms is labeled by $d_{1}$ in Fig. 1(a), and the values are listed in Table 1. The general trend is that $d_{1}$ and $d_{2}$ values increase with heavier ions in G4V(--) centers at zero pressure. Such behavior may be expected as a larger ion creates a stronger strain field around the dopant ion. For a given G4V(--) center, the $d_{1}$ and $d_{2}$ values slowly decrease with the increase of hydrostatic pressure. This response may be also expected as the hydrostatic pressure generally decrease the bond length in the diamond crystal which also holds for the bond lengths in the center of G4V defects.
@@
SiV(--) 1.97///1.96111Ref. 26 2.67///2.67111Ref. 26 1.79 2.40
GeV(--) 2.01///2.01111Ref. 26 2.74///2.73111Ref. 26 1.84 2.48
SnV(--) 2.09///2.08111Ref. 26 2.83///2.83111Ref. 26 1.92 2.58
PbV(--) 2.12///2.12111Ref. 26 2.89///2.88111Ref. 26 1.96 2.64
@@
Table 2: The calculated Jahn-Teller (JT) energy in meV ($E_{\text{JT}}$), the barrier energy in meV ($\delta_{\text{JT}}$), the effective mode in meV ($\hbar\omega$), spin-orbit splitting $\lambda_{0}$ (meV), $p$ reduction factor, $g$ the orbital reduction factor, and $\lambda$ (GHz) which is Ham reduced $\lambda_{0}$ for the G4V(--) defects ($\lambda=p\lambda_{0}$) with no applied pressure. The left/right hand side data are from SCAN/HSE06 calculations.
Defects $E_{\text{JT}}$ (meV) $\delta_{\text{JT}}$ (meV) $\hbar\omega$ (meV) $\lambda_{\text{0}}$ (meV) $p$ $g$ $\lambda$ (GHz) $\lambda_{\text{exp}}$ (GHz)
@@
SiV(--) 40.9 / 42.3111Ref. 26 3.79 / 3.0111Ref. 26 89.7 / 85.2111Ref. 26 0.86 / 0.82111Ref. 26 0.34 / 0.31111Ref. 26 0.328111Ref. 26 70.26 / 61.0111Ref. 26 50222Ref. 2
GeV(--) 30.6 / 30.1111Ref. 26 4.05 / 2.0111Ref. 26 77.0 / 82.2111Ref. 26 2.45 / 2.20111Ref. 26 0.38 / 0.39111Ref. 26 0.328111Ref. 26 222.8 / 207111Ref. 26 181333Ref. 5
SnV(--) 20.8 / 21.6111Ref. 26 1.15 / 1.6111Ref. 26 64.9 / 79.4111Ref. 26 8.69 / 8.28111Ref. 26 0.44 / 0.47111Ref. 26 0.328111Ref. 26 915.2 / 946111Ref. 26 850444Ref. 8
PbV(--) 15.0 / 15.6111Ref. 26 3.88 / 0.6111Ref. 26 52.0 / 74.9111Ref. 26 35.0 / 34.6111Ref. 26 0.48 / 0.54111Ref. 26 0.328111Ref. 26 4097 / 4514111Ref. 26 3914555Ref. 18
@@
SiV(--) 62.6 / 78.5111Ref. 26 1.12 / 2.7111Ref. 26 61.0 / 73.5111Ref. 26 8.864 / 6.96111Ref. 26 0.133 / 0.128111Ref. 26 0.782111Ref. 26 286.2 / 215111Ref. 26 260222Ref. 2
GeV(--) 71.5 / 85.7111Ref. 26 2.31 / 5.4111Ref. 26 70.6 / 73.0111Ref. 26 35.03 / 36.1111Ref. 26 0.136 / 0.113111Ref. 26 0.782111Ref. 26 1155 / 987111Ref. 26 1120333Ref. 5
SnV(--) 67.7 / 83.1111Ref. 26 4.20 / 6.8111Ref. 26 68.1 / 75.6111Ref. 26 94.77 / 96.8111Ref. 26 0.140 / 0.125111Ref. 26 0.782111Ref. 26 3214 / 2897111Ref. 26 3000444Ref. 8
PbV(--) 87.3 / 91.6111Ref. 26 6.69 / 12.3111Ref. 26 77.9 / 78.6111Ref. 26 241.4 / 245111Ref. 26 0.116 / 0.119111Ref. 26 0.782111Ref. 26 6782 / 7051111Ref. 26 -
For each applied hydrostatic pressure, we calculated the APES by SCAN functional, as illustrated in Fig. 1(d). For the zero-pressure case, we compared the SCAN data with previous HSE06 data (see Table 2). We found a good agreement between the SCAN and HSE06 data for the critical JT parameters (such as the Jahn–Teller energy $E_{\text{JT}}$, barrier energy $\delta_{\text{JT}}$, and the effective frequency $\hbar\omega$), so we continued the calculations with SCAN functional for non-zero pressure cases. Table 9 in the Appendix lists the respective JT parameters for G4V(--) centers under pressure.
@@ III.3 Effective spin-orbit splitting
We calculated the spin-orbit coupling using the SCAN functional, which is an essential parameter in capturing the fine electronic structure of G4V(--) centers. We label this electronic spin-orbit coupling as $\lambda_{0}$, listed for each G4V(--) for the zero-pressure case in Table 2. Our results show strong agreement between the SCAN data and the previously published HSE06 $\lambda_{0}$ values for each electronic state in the G4V(--) color centers. This consistency supports confidence that using SCAN functional, the electronic spin-orbit coupling can be further calculated for the G4V(--) color centers under non-zero hydrostatic pressures.
@@
Figure 3: Pressure dependent spin-orbit splitting as obtained with SCAN functional where the spin-orbit interaction added as perturbation to the Jahn–Teller-effect. (a) The calculated hydrostatic pressure (GPa) dependence of the effective spin-orbit splitting $\lambda$ in GHz unit for G4V(--) color centers in the electronic excited and ground states plotted by blue and red lines, respectively. (b) The sum of the electronic excited and ground state’s $\lambda$ values associated with the zero-phonon line broadening at elevated measurement temperatures with off-resonant excitation for each G4V(--) color center as a function of the applied hydrostatic pressure.
@@
In Jahn-Teller active systems, the electronic spin-orbit coupling is reduced due to strong electron-phonon coupling, and this reduction is quantified by the Ham reduction factor that we label by $p$. In our previous work [62], we described the ab initio calculation of the Ham reduction factor for the $E\otimes e$ Jahn-Teller system, which was later applied to G4V(--) defects in a subsequent study [26]. For the zero pressure case, we find that the calculated $p$ factors and the resulting $\lambda$ for the electronic ground and excited state of G4V(--), as obtained by SCAN and HSE06 functionals, are consistent. The calculated $\lambda$ values also show good agreement with experimental data for both functionals (see Table 2). This result further supports the applicability of the SCAN functional for investigating the dependence of effective spin-orbit splitting on hydrostatic pressure in G4V(--) color centers.
@@
Fig. 3(a) shows the calculated effective spin-orbit splitting as a function of hydrostatic pressure, obtained using SCAN functional. The general trend is that the effective spin-orbit splitting is monotonously increasing with increasing hydrostatic pressure both in the electronic ground and excited states. The calculated effective spin-orbit splitting has two factors: the Ham-reduction parameter $p$ and the calculated electronic spin-orbit splitting $\lambda_{0}$ (see Table 9). We find that the Ham-reduction parameter only slightly varies as a function of hydrostatic pressure. Therefore, the primary contribution to the overall increase in spin-orbit coupling under rising hydrostatic pressure comes from the inherent electronic spin-orbit splitting. This change can be explained by analyzing the dominant contribution to the electronic spin-orbit splitting, which arises from the spin-orbit coupling on the dopant ion [26]. As hydrostatic pressure increases, the distance ($d_{1}$) between the dopant ion and the neighbor carbon atoms decreases, enhancing the overlap of the dangling bonds with the dopant ion (see also Sec. III.4). Consequently, the wavefunction cloud around the dopant ion becomes denser, contributing significantly to the spin-orbit coupling. The steepest shift in the effective spin-orbit splitting is observed for the SnV(--) color center.
@@
In Fig. 3(b), we plot the sum of $\lambda_{g}$ and $\lambda_{u}$ – defined in Fig. 1(c) – as a function of the applied hydrostatic pressure, which we discuss in the context of the broadening of the ZPL emission. Off-resonant excitation of G4V(--) color centers should result in emission in all the possible combinations of states and levels depicted in Fig. 1(c). By assuming a homogeneous strain field and neglecting the isotope effects (in the GHz region), the sum of $\lambda_{g}$ and $\lambda_{u}$ yields the width of the zero-phonon line. In experiments, the broadening of ZPL for SiV(--) varied from 10 to 30 meV going from zero to $180$ GPa of hydrostatic pressure, whereas it varied from 15 to 30 meV for GeV(--), respectively [19]. In our calculations, the spin-orbit splitting related broadening of the ZPL emission varies between 1.4 meV to 1.8 meV for SiV(--) going from zero to $180$ GPa of hydrostatic pressure, whereas it varies between 5.5 meV to 6.7 meV for GeV(--), respectively. The calculated values are smaller in order of magnitude both in terms of absolute values and changes upon hydrostatic pressures. We conclude that before applying pressure to the sample, non-uniform local strain fields could be already present in the samples, which causes inhomogeneous broadening in the region of 10 meV. Applying hydrostatic pressure to the samples could further increase the variety of the local strain fields that may explain the observed wide broadening of the ZPL emission for SiV(--) and GeV(--) optical centers.